OpenStudy (anonymous):

Ok so i think i have the answers but i think its wrong so please help! Btw this is so be solved using Combination Notation (nCr): The letters A, B, C,D, E, F,G, H, I, and J are written on slips of paper and placed in a hat.Two letters are then drawn from the hat. a. What is the number of possible combinations of two letters? b. How many combinations consist only of the vowels A, E, or I? c. What is the probability that the letters chosen would consist only of vowels? d. What is the probability that the letters chosen would be only B, C,D, or F?

OpenStudy (anonymous):

ten letter answer to number 1 is 10*9/2=45

OpenStudy (anonymous):

10C2

OpenStudy (anonymous):

b there are 3

OpenStudy (anonymous):

@satellite73 thats what i got :)

OpenStudy (anonymous):

c is 3 out of 45

OpenStudy (anonymous):

a. 10*9=90 b. 3*2=6 c. 3*2/10*9=1/15 d. 4*3/10*9=2/15

OpenStudy (anonymous):

90 is not right. have to divide by 2

OpenStudy (anonymous):

a is not 10*9.

OpenStudy (anonymous):

@rnoy i got those for c and d but not for a and b.

OpenStudy (anonymous):

d 4 choose 2 is 4*3/2=6

OpenStudy (anonymous):

so 6/45

OpenStudy (anonymous):

b is 3C2 = 3

OpenStudy (anonymous):

for a i got 3,6238,800 b. 3 c. 1/15 d. 2/15

OpenStudy (anonymous):

a is not right.

OpenStudy (anonymous):

10C2 = 45

OpenStudy (anonymous):

@polpak i think it is... i need the total number of outcomes! for A its 9 then B 8 the C 7... i think

OpenStudy (anonymous):

No, from the set of 10 letters you are choosing 2. There are only 45 different combinations.

OpenStudy (anonymous):

There are only 45 different ways you can pick 2 things from a set of 10. 10C2

OpenStudy (clover):

thats what i got

OpenStudy (anonymous):

ab ac ad ae af ag ah ai aj thats 9 and then for b 8 and c 7 and d 6 and so on

OpenStudy (anonymous):

@polpak the question asks not only to choose but, the combination is also asked (eg. difference between [A,B],[B,A] should be considered)

OpenStudy (anonymous):

Except you are not removing duplicates.

OpenStudy (anonymous):

And even permutations only gives 90 different possibilities.

OpenStudy (anonymous):

@polapak i did!

OpenStudy (clover):

they did

OpenStudy (anonymous):

ok well a isnt important! well not too much b c and d are a little more important

OpenStudy (anonymous):

9 + 8 + 7 + 6 + 5 + 4 + 3 + 2 + 1 = 45 = 10C2

OpenStudy (anonymous):

its not plus!!!! its times! clover isnt that right?

OpenStudy (clover):

thats wrong

OpenStudy (anonymous):

b is just 3C2. (picking 2 elements from a group of 3).

OpenStudy (clover):

kas9987 is right....its times

OpenStudy (anonymous):

but isnt is ae ai ei?

OpenStudy (anonymous):

No it's not. You were enumerating all the different combinations. If you want to know how many there are you add 1 for each combination.

OpenStudy (anonymous):

oh sorrry that equals the same thing!

OpenStudy (anonymous):

\[ei \equiv ie\]

OpenStudy (anonymous):

That's why it's call 'choose'. If you want to know how many ways you can possibly pick k items from a set of n it's nCk

OpenStudy (clover):

you mean nCr

OpenStudy (anonymous):

no. I mean what I said.

OpenStudy (anonymous):

i think so but 3C2 = (3*2)/(2*1) which is 6/2=3

OpenStudy (anonymous):

Yes. 3 is the correct answer for b, and 45 is the correct answer for a.

OpenStudy (clover):

im pretty sure its nCr

OpenStudy (anonymous):

And what is r?

OpenStudy (anonymous):

no its nCr! We just learned this today! just wondering what type of are you/ have you taken/ passed?

OpenStudy (anonymous):

I'm just about finished with this nonsense. I've taken discrete math, linear algebra, multi variable calculus, and differential equations. If you want to call it nCr that's fine, but you need to read what I wrote and realize that what you are calling 'r' (the number of items to pick) I called k and it doesn't matter if I called it argleblargle it would mean the same thing.

OpenStudy (anonymous):

nCargleblargle is the number of ways you can possibly pick argleblargle items from a set of n items.

OpenStudy (anonymous):

ok! we're sorry!

OpenStudy (clover):

hi

OpenStudy (anonymous):

Ok, so now the question is.. How many items do we have to pick from?

OpenStudy (anonymous):

(if we're talking about part a)

OpenStudy (anonymous):

10 i think... A-j

OpenStudy (anonymous):

Right. And how many items are we picking from that set of 10?

OpenStudy (anonymous):

2 but one u pick one u only have 9 to choose from... wait! i think u may be right... I'm really sorry! thank you!

OpenStudy (anonymous):

I mean we can list out all the combinations if you want.. it'll take a while, but it's not impossible.

OpenStudy (anonymous):

im sorry im just a butt head! i was thinking of it was though i had to pick all of the letters out... I'm really sorry!

OpenStudy (anonymous):

First all the ones with a: ab,ac,ad,ae,af,ag,ah,ai,aj Then all the ones with b (minus the ones with a): bc,bd,be,bf,bg,bg,bi,bj Then all the ones with c (minus any with a or b): cd,ce,cf,cg,ch,ci,cj etc. de,df,dg,dh,di,dj etc. ef,eg,eh,ei,ej etc. fg,fh,fi,fj etc. gh,gi,gj etc. hi,hj etc. ij 1 + 2 + 3 + 4+ 5+ 6 + 7 + 8 + 9 = 10C2 = 45

OpenStudy (anonymous):

For part c, how many vowels are there?

OpenStudy (anonymous):

really im so sorry! thank you so much for being patient! and a, e and i so 3!

OpenStudy (anonymous):

And so how many different possible ways are there to choose 2 from that set of vowels?

OpenStudy (anonymous):

hint: you already answered this question.

OpenStudy (anonymous):

hi polpak. did you convince them is was \[\dbinom{n}{k}\] and not permutations yet?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

whew!

OpenStudy (anonymous):

Kas, there are 3C2 = 3 ways to pick 2 letters such that they are only vowels. But there are 10C2 = 45 different possible ways you can pick 2 letters from the 10 you have. So the probability will be \[{3C2\over 10C2} = {3 \over 45} = {1 \over 15}\]

OpenStudy (anonymous):

For part d the logic will be similar but you have 4C2 different combinations you can make with the letters listed so you have: \[{4C2 \over 10C2} = {6 \over 45} = {1 \over 7}\]

OpenStudy (anonymous):

umm thank you so much but one thing... 6/45 simplified is 2/15 no 1/7... i just wanted to point that out but thank you so much! :)